The sum of the ages of Brain,steve and phil is 53.brain's age now is twice steve's age five years ago,and brain is two years older than Phil.What is the age of the oldest? Please show me the steps.
A.19 B.25 C.20 D.18 E.None of these

Answer by flame8855(424)

let b be the age of brain
let s be the age of steve
let p be the age of phil
b+s+p=53
b=2(s-5)
s=(b+10)/2
b=p-2
p=b+2
substitute
b+(b+10)/2 +b+2 = 53
2b+b+10+2b+4=106
5b= 92
b=92/5=18.4
s=(b+10)/2=14.2
p=b+2= 20.4
so the age of the oldest is 20